Find values for m and b in the following system so that the solution to the system is ( 1, 2).
mx+ 3y = 8
-3x + 4y = b

Answer by malakumar_kos@yahoo.com(315)


Find values for m and b in the following system so that the solution to the system is ( 1, 2).
mx+ 3y = 8
-3x + 4y = b
Given eq'n is mx+3y = 8.....(1)
-3x+4y = b.....(2)
substitute for x,y in both the eq'ns
m*1+3*2 = 8 or m+6 = 8 or m = 8-6 = 2
-3(1)+4(2) = b or -3+8 = b or b= 8-3 = 5
Therefore the values of m & b are 2 & 5